OpenStudy (anonymous):

Solve the differential equation y"+y=(secx)^2.

OpenStudy (amistre64):

r^2 + 1 = 0 (r+i) (r-i) r = complex roots r= (0+i1, 0-i1) yh= e^0 (C1 cos(x)+C2 sin(x)) yh= C1 cos(x) + C2 sin(x)

OpenStudy (amistre64):

yp = A cos(x) + Bsin(x) yp' = -Asin(x) + Bcos(x) +A'cos(x) + B'sin(x) --------------- = 0 yp' = -Asin(x) + Bcos(x) yp'' = -Acos(x) - Bsin(x) -A'sin(x) + B'cos(x) + yp = A cos(x) + Bsin(x) --------------------------------------- sec^2(x) = -A'sin(x) + B'cos(x)

OpenStudy (amistre64):

A'cos(x) + B'sin(x) = 0 -A'sin(x) + B'cos(x) = sec^2(x) system of equation, solve for A' and B'

OpenStudy (amistre64):

If we cramer it B' = sec^2 cos / cos^2 + sin^2 = sec A' = sec^2 sin / -sin^2 - cos^2 = -tan sin integrate A' and B' to find A and B and sub them into yp the solution then is: y = yh + yp

OpenStudy (anonymous):

in the book they say that something about W.....

OpenStudy (amistre64):

Whats George Bush got to do with this?